OpenStudy (anonymous):

25 increased by 200%

OpenStudy (anonymous):

200% of 25 = (200/100)* 25 =50 so the answer is 25+50 = 75

OpenStudy (anonymous):

can you help me with another one?

OpenStudy (anonymous):

sure